OpenStudy (anonymous):

guys i really need your help with this one, each time I get a different answer and I don't know what im doing wrong I need you geniuses to help me...pretty please.... solve by substitution 15x-20y=1 15y=-2+5x

OpenStudy (experimentx):

first seperate x from first equation, then put the value of x in the second equation so that you have only y, and solve y, then find the value of x from any equation

OpenStudy (anonymous):

its all crazy e, i get all kinds of fractions and decimals ok look

OpenStudy (anonymous):

y=1-15xover 20

OpenStudy (anonymous):

y=1+5/4x

OpenStudy (experimentx):

x = (1+20y)/15 and then 15y = -2+5*(1+20y)/15 => y = -1/5 and so on

OpenStudy (anonymous):

no y=1+3/4x

OpenStudy (anonymous):

ok im gonna try but I do think I got that answer already yup -1/5 let me rerererertry it lol